### v1

@@ 3.5 Model estimation via multiple shooting and subspace encoding
To overcome problem (ii), we truncate the $\ell_{2}$ prediction loss (33) to a horizon of length $T$ and we divide the data into subsections on which the truncated prediction loss is calculated, giving a so called multiple shooting form of the optimization problem. This approach reduces the computational cost and improves numerical stability [37]. Accordingly, the prediction loss is reformulated as:
@@
[…]
which enables the application of advanced batch optimization algorithms like Adam [18]. Moreover, the complete dataset does not need to be fully loaded into the memory, making the implementation more efficient [2].
@@ 4 Consistency analysis
Next, we show the consistency of the proposed identification scheme that corresponds to the minimization of (38). In fact, under the assumption that an exact finite-dimensional Koopman embedding of (21) in the form of (25) exists, we show that the resulting model estimate will converge to an equivalent representation of the system in the form (25) if the number of data points in the available data set $\mathcal{D}_{N}$ tends to infinity, that is, $N\rightarrow\infty$. The consistency analysis discussed in this section is an adaptation of the arguments in [2] to the considered Koopman identification problem.
@@ 4.1 Convergence
To identify (25), the model structure $M_{\xi}$ is composed of the Koopman model (32) with the forms of parametrization discussed in Section 3.3 and the subspace encoder (39) with the parametrization discussed in Section 3.4, giving the total parameter vector $\xi=[\begin{array}[]{cc}\theta^{\top}&\eta^{\top}\end{array}]^{\top}$ that is restricted to vary in a compact set $\Xi\subset\mathbb{R}^{n_{\xi}}$. This gives the model set =$\mathcal{M}=\{M_{\xi}\mid\xi\in\Xi\}$. For each $\xi\in\Xi$, the model $M_{\xi}$ with a given encoder lag $n\geq 1$, can be written in a one-step-ahead predictor form
@@
If the Koopman form (25) of the data-generating system satisfies Condition 1 with a quasi-stationary $u$ independent of the white noise process ${e}$ and the model set $\mathcal{M}$ defined by (32) and (39) satisfies Condition 2, then
@@
As the identification criterion (38a) satisfies Condition C1 in [21], the proof of [21, Lemma 3.1] applies to the case considered.
@@ 4.2 Consistency
For the given model set =$\mathcal{M}=\{M_{\xi}\mid\xi\in\Xi\}$ with $\xi=[\begin{array}[]{cc}\theta^{\top}&\eta^{\top}\end{array}]^{\top}$, we call the input sequence $\{u_{k}\}_{k=0}^{N-1}$ in $\mathcal{D}_{N}$ generated by the Koopman form (25) of the data-generating system weakly persistently exciting, if for all pairs of parameters given by $\xi_{1}\in\Xi$ and $\xi_{2}\in\Xi$ for which the function mapping is unequal, i.e., $V_{(\cdot)}^{\mathrm{enc}}(\theta_{1},\eta_{1})\neq V_{(\cdot)}^{\mathrm{enc}}(\theta_{2},\eta_{2})$, we have
@@
Next, to prove consistency, all elements of $\Xi_{\mathrm{o}}$ must have minimal asymptotic cost in terms of $\lim_{N\rightarrow\infty}V_{\mathcal{D}_{N}}^{\mathrm{enc}}(\theta,\eta)$. However, due to the prediction error nature of the used $\ell_{2}$-type loss function together with the existence of $\mathbb{E}_{e}\{V_{\mathcal{D}_{N}}^{\mathrm{end}}(\theta,\eta)\}$ (shown in Theorem 4.1), the minimal asymptotic cost property of $\Xi_{\mathrm{o}}$ is satisfied by $V_{\mathcal{D}_{N}}^{\mathrm{enc}}$. For a detailed proof, see [21].
@@
The proof is a direct application of Lemma 4.1 in [21] because the loss function (38a) fulfills Condition (4.4) in [21].
@@ 5.1 Wiener-Hammerstein system
with $\bar{x}_{k}\in\mathbb{R}^{n_{\mathrm{\bar{x}}}}$ the state vector and with matrices similarly defined as for the first LTI block. With $n_{\mathrm{x}}=n_{\bar{\mathrm{x}}}=2$, the exact numerical values of the matrices and the polynomial coefficients are given in [13]. The system described by (48)–(51) can be exactly represented by a finite dimensional Koopman model (25). For brevity, we refer the reader to [13] for the derivations, which uses a similar finite dimensional conversion approach to [11]. The resulting lifted state is
@@
To generate data, the input is considered as a white noise process with uniform distribution $u_{k}\sim\mathcal{U}(-1,1)$, independnent of $e$, while the standard deviation $\sigma_{e}$ of $e$ is chosen to obtain $5-30$ dB levels of signal-to-noise ratio (SNR) at the output. Based on this, train, validation and test data sets are generated of length $N=12000$, 4000 and 4000, respectively, with independent realizations of $u$ and $e$.
@@
In the considered Koopman model structure $M_{\xi}$, defined by (32) and (39), the encoder $\bar{\Pi}^{\eta}_{\mathrm{z},n}$, the input function $B_{\theta}$ and the innovation noise structure $K_{\theta}$ are parametrized as feedforward neural networks with 1 hidden layer, tanh activation and 40 neurons per layer for the encoder and $B_{\theta}$ function, and 80 neurons for $K_{\theta}$, while $A_{\theta}$ and $C_{\theta}$ are taken as parametrized matrices (also $K_{\theta}$ in the linear case). The parameters are initialized using Xavier initialization and we employ early stopping. The lifting dimension is selected to coincide with the exact model, i.e., $n_{z}=12$ and we use an encoder lag of $n=12$. The prediction horizon is chosen to be $T=51$ with a batch size of 256. For training, Adam optimization [18] is used with the obtained training and validation data sets and with a learning rate of $\alpha=10^{-3}$ and the exponential decay rates set to $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The quality of the obtained models are assessed in terms of the normalized root mean square (NRMS) and RMS errors:
@@
Fig. 4 shows the simulation performance of the identified models on noiseless test data when trained on noisy data of particular SNRs. Models with various choices of $K_{\theta}$ are also compared: no $K_{\theta}$, linear $K_{\theta}$, and general $K_{\theta}(\hat{z}_{k},u_{k},\hat{e}_{k})$. It can be seen that the linear innovation noise structure is able to reduce the approximation error with a factor of up to two and slight improvements are obtained with the more complex noise model. Next, we analyse the one-step-ahead prediction performance. Similar to the simulation test case, Fig. 5 compares the NRMS errors of these model structures, in a filtering scenario, using noisy test data. As in the simulation scenario, the models are trained on noisy data with the respective SNRs. To provide context for these results, two additional dashed reference lines are included in the figure which represent the one-step-ahead prediction error of the true model, showing the proven convergence and consistency properties as the noise diminishes. It is also clear that incorporating a linear $K_{\theta}$ significantly reduces the NRMS error compared to the baseline model (which is correctly identified) without an innovation noise structure. Overall, both the simulation and prediction results highlight the importance of the innovation noise structure in increasing accuracy of the identified Koopman models.
@@ 5.2 Bouc-Wen benchmark
The Bouc-Wen oscillator benchmark [32, 33, 40] is extensively used for testing capabilities of nonlinear system identification approaches as it describes a system with hysteresis which is a challenging behavior to capture from data accurately. The Bouc-Wen oscillator can be modeled as:
@@
with $m_{L}$ the mass, $y$ its displacement, $\dot{y}$ the velocity, and $u$ the external force applied to the system. The restoring force $r(y,\dot{y})$ is linear while $q(y,\dot{y})$ is a dynamic nonlinear function describing the hysteresis curve. Both these effects are extensively described in [40] where the parameters are chosen for the Bouc-Wen benchmark such that the hysteretic behavior is significantly present.
To estimate a Koopman model of the Bouc-Wen benchmark system, the training dataset contains input-output data generated by using two sinusoidal signals with frequencies of 1 and 4 Hz, as well as four random phase multisine signals with frequency bands ranging from 1 to 150 Hz. For validation, a dataset is generated with an input containing a multisine signal also with an excited frequency range between 1 and 150 Hz followed by a sinusoidal signal at 2 Hz. For test data, the simulated response is generated by multisine and sinesweeep signals as given in [40], as well as a 3 Hz sinusoidal signal to showcase the hysteretic behavior. Note that, as detailed in [40], the data is sampled at a frequency of $f_{\mathrm{s}}=750$ Hz. These used detasets are shown in Fig. 6. As the data is noiseless, the estimation of a noise model is not required which means that $K_{\theta}$ is set to 00.
@@
The best performing model in the multisine case is the model structure with input affine complexity and a lifting dimension of $n_{\mathrm{z}}=100$ (although a close result is also obtained with $n_{\mathrm{z}}=20$). We use this model for comparing the results against other available methods. In Fig. 8, the simulation results using the test data show a low error for the multisine and sinesweep input excitation, and clearly show that the hysteretic behavior is well captured. In Table 1, we can see that the simulation performance of the obtained Koopman model is close to the state of the art (for the interested reader, the other methods are described in more detail in [8, 39]). The approaches that obtain slightly better results do not impose a particular structure on the learned model. The Koopman model (25) is able to accurately capture the system behavior and offers a good overall trade-off between state dimension and model complexity. Furthermore, even the more complex structures, i.e., input affine or general, can be cast into LPV representations [12] for which there exist convex and computationally efficient control methods [30].
Method gr-SS-NN SS-NN Suykens Impr SS-NN Suykens Poly-NL-SS
@@ 5.3.1 Simulation study
The considered simulation model of the drone implements the rigid body dynamics as described in [25] and uses three coordinate frames: north-east-down (NED) oriented inertial frame $F_{\mathrm{i}}$; the vehicle frame $F_{\mathrm{v}}$ (origin at the centre of gravity of the quadrotor) sharing the same orientation as $F_{\mathrm{i}}$ ; the body frame $F_{\mathrm{b}}$ (orientation fixed to the quadrotor) whose origin coincides with $F_{\mathrm{v}}$. The model has 12 motion states composed of the position $s=[\ x\ \ y\ \ z\ ]^{\top}$, translational velocity $v=[\ v_{\mathrm{x}}\ \ v_{\mathrm{y}}\ \ v_{\mathrm{z}}\ ]^{\top}$ expressed in $F_{\mathrm{i}}$, $\zeta=[\ \phi\ \ \theta\ \ \psi\ ]^{\top}$ describing the orientation as Z-Y-X Euler angles in $F_{\mathrm{v}}$, and $\omega=[\ p\ \ q\ \ r]^{\top}$, describing the rotational velocity of $F_{\mathrm{b}}$ w.r.t. $F_{\mathrm{v}}$, given in $F_{\mathrm{b}}$. The inputs to the system are the total thrust $T$ and the torque vector $\tau=[\ \tau_{\phi}\ \ \tau_{\theta}\ \ \tau_{\psi}]^{\top}$ both given in $F_{\mathrm{b}}$ and produced by the four rotors. The diagonal values of the inertia matrix of the drone are set to $J_{\mathrm{x}}=J_{\mathrm{y}}=1.4\times 10^{-5}\,\text{kg}\!\cdot\!\text{m}^{2}$, $J_{\mathrm{z}}=2.17\times 10^{-5}\,\text{kg}\!\cdot\!\text{m}^{2}$, and the off-diagonal values are 0, the mass is $m=0.027\,\text{kg}$, and $g=9.8\,\text{m/s}^{2}$, which are experimentally obtained using a real-world Crazyflie 2.1 quadrotor. The simulation is performed with Runge-Kutta 4 integration at a sampling rate of 48 Hz, accurately replicating the expected flight-dynamics.
As the system is unstable, flight-trajectories are generated by using a gain-scheduled linear quadratic regulator (LQR) controller, designed w.r.t. the local linearisations at each time step of the simulation model. The LQR is scheduled based on a desired state trajectory that is calculated for a $x$-$y$-$z$-$\psi$ defined flight-path reference by taking advantage of the differential flatness property of the system, see [29]. We presume full state measurements, hence the collected dataset consists of the system inputs and the states of the dynamical system. Since the dynamics governing the evolution of the position states consist solely of integration, the identification process can be simplified by focusing only on $[\ v^{\top}\ \ \zeta^{\top}\ \ \omega^{\top}\ ]^{\top}$. The size of the recorded datasets can be viewed in Table 10. For more details about the simulation and data collection procedure, the reader can refer to [47].
[…]
@@
Model performance across various $n_{\mathrm{z}}$ values can be seen in Table 2 in terms of the 160-step NRMS errors for the trained models on the test data set. For comparison purposes, results on a full nonlinear model estimated by the SUBNET approach [47] are also included to indicate performance of a nonlinear model estimate without any restrictions on the network structure. In the network structure column, the subscripts $(\cdot)_{\theta}$, $(\cdot)_{\text{lin}}$, and $(\cdot)_{\mathbf{I}}$ denote the implementation of a function as a deep neural network, linear or identity layers. The superscript of $A^{(\cdot)}$ denotes the lifted state-space dimension. The Koopman model achieves good performance in terms of the NRMS error only slightly exceeding that of the nonlinear SUBNET. Similar to the Bouc-Wen example, we found that the dependence of $B$ on $u$ slightly deteriorates the performance, compared to the input affine model structure. In the table, the effects of measurement noise can also be seen. Even at a level of 20 dB, the Koopman model remains capable of capturing the dynamics. Enforcing the original states among the observables only slightly decreases the simulation precision of the model, which we consider as a good tradeoff for the simpler model structure.
@@
Simulation $A^{20}$, $B_{\theta}(z)$, $C_{\text{lin}}$ None 0.108 0.120 0.130
@@
Real $A^{40}$, $B_{\theta}(z)$, $C_{\text{lin}}$ Sensor [^3] No artificial noise was added, the real dataset was only affected by the noise inherent to the inaccuracies of the various sensors equipped on the Crazyflie. 0.163 0.213 0.228

### v3

@@ 3.5 Model estimation via multiple shooting and subspace encoding
To overcome problem (ii), we truncate the $\ell_{2}$ prediction loss (33) to a horizon of length $T$ and we divide the data into subsections on which the truncated prediction loss is calculated, giving a so called multiple shooting form of the optimization problem. This approach reduces the computational cost and improves numerical stability [Ribeiro]. Accordingly, the prediction loss is reformulated as:
@@
(40c) $\text{s.t. }|\mathcal{I}|=N_{\text{batch}}$
which enables the application of advanced batch optimization algorithms like Adam [adam]. Moreover, the complete dataset does not need to be fully loaded into the memory, making the implementation more efficient [Gerben_Aut].
@@ 4 Consistency analysis
Next, we show the consistency of the proposed identification scheme that corresponds to the minimization of (38). In fact, under the assumption that an exact finite-dimensional Koopman embedding of (21) in the form of (25) exists, we show that the resulting model estimate will converge to an equivalent representation of the system in the form (25) if the number of data points in the available data set $\mathcal{D}_{N}$ tends to infinity, that is, $N\rightarrow\infty$. The consistency analysis discussed in this section is an adaptation of the arguments in [Gerben_Aut] to the considered Koopman identification problem.
@@ 4.1 Convergence
To identify (25), the model structure $M_{\xi}$ is composed of the Koopman model (32) with the forms of parametrization discussed in Section 3.3 and the subspace encoder (39) with the parametrization discussed in Section 3.4, giving the total parameter vector $\xi=[\begin{array}[]{cc}\theta^{\top}&\eta^{\top}\end{array}]^{\top}$ that is restricted to vary in a compact set $\Xi\subset\mathbb{R}^{n_{\xi}}$. This gives the model set $\mathcal{M}=\{M_{\xi}\mid\xi\in\Xi\}$. For each $\xi\in\Xi$, the model $M_{\xi}$ with a given encoder lag $n\geq 1$, can be written in a one-step-ahead predictor form
@@
If the Koopman form (25) of the data-generating system satisfies Condition 1 with a quasi-stationary $u$ independent of the white noise process e{e} and the model set $\mathcal{M}$ defined by (32) and (39) satisfies Condition 2, then
@@
As the identification criterion (38a) satisfies Condition C1 in [ljung1978convergence], the proof of [ljung1978convergence, Lemma 3.1] applies to the case considered.
@@ 4.2 Consistency
For the given model set $\mathcal{M}=\{M_{\xi}\mid\xi\in\Xi\}$ with $\xi=[\begin{array}[]{cc}\theta^{\top}&\eta^{\top}\end{array}]^{\top}$, we call the input sequence $\{u_{k}\}_{k=0}^{N-1}$ in $\mathcal{D}_{N}$ generated by the Koopman form (25) of the data-generating system weakly persistently exciting [^5] Note that the notion of persistency of excitation used here is in line with the classical notion of informativity in system identification, see [book_ljung], and it implies distinguishability (under the data $\mathcal{D}_{N}$) of the achieved cost and the predictors corresponding to the used model structure for $\theta$ values that do not correspond to equivalent models., if for all pairs of parameters given by $\xi_{1}\in\Xi$ and $\xi_{2}\in\Xi$ for which the function mapping is unequal, i.e., $V_{(\cdot)}^{\mathrm{enc}}(\theta_{1},\eta_{1})\neq V_{(\cdot)}^{\mathrm{enc}}(\theta_{2},\eta_{2})$, we have
@@
Next, to prove consistency, all elements of $\Xi_{\mathrm{o}}$ must have minimal asymptotic cost in terms of $\lim_{N\rightarrow\infty}V_{\mathcal{D}_{N}}^{\mathrm{enc}}(\theta,\eta)$. However, due to the prediction error nature of the used $\ell_{2}$-type loss function together with the existence of $\mathbb{E}_{e}\{V_{\mathcal{D}_{N}}^{\mathrm{end}}(\theta,\eta)\}$ (shown in Theorem 4.1), the minimal asymptotic cost property of $\Xi_{\mathrm{o}}$ is satisfied by $V_{\mathcal{D}_{N}}^{\mathrm{enc}}$. For a detailed proof, see [ljung1978convergence].
@@
The proof is a direct application of Lemma 4.1 in [ljung1978convergence] because the loss function (38a) fulfills Condition (4.4) in [ljung1978convergence].
@@ 5.1 Wiener-Hammerstein system
with $\bar{x}_{k}\in\mathbb{R}^{n_{\mathrm{\bar{x}}}}$ the state vector and with matrices similarly defined as for the first LTI block. With $n_{\mathrm{x}}=n_{\bar{\mathrm{x}}}=2$, the exact numerical values of the matrices and the polynomial coefficients are given in [Iacob_DT_WH_embedding]. The system described by (48)–(51) can be exactly represented by a finite dimensional Koopman model (25). For brevity, we refer the reader to [Iacob_DT_WH_embedding] for the derivations, which uses a similar finite dimensional conversion approach to [Iacob_CT_block_embedding]. The resulting lifted state is
@@
To generate data, the input is considered as a white noise process with uniform distribution $u_{k}\sim\mathcal{U}(-1,1)$, independent of $e$, while the standard deviation $\sigma_{e}$ of $e$ is chosen to obtain 5305-30 dB levels of signal-to-noise ratio (SNR) at the output. Based on this, train, validation and test data sets are generated of length $N=12000$, 4000 and 4000, respectively, with independent realizations of $u$ and $e$.
@@
In the considered Koopman model structure $M_{\xi}$, defined by (32) and (39), the encoder $\bar{\Pi}^{\eta}_{\mathrm{z},n}$, the input function $B_{\theta}$ and the innovation noise structure $K_{\theta}$ are parametrized as feedforward neural networks with 1 hidden layer, tanh activation and 40 neurons per layer for the encoder and $B_{\theta}$ function, and 80 neurons for $K_{\theta}$, while $A_{\theta}$ and $C_{\theta}$ are taken as parametrized matrices (also $K_{\theta}$ in the linear case). The parameters are initialized using Xavier initialization and we employ early stopping. The lifting dimension is selected to coincide with the exact model, i.e., $n_{z}=12$ and we use an encoder lag of $n=12$. The prediction horizon is chosen to be $T=51$ with a batch size of 256. For training, Adam optimization [adam] is used with the obtained training and validation data sets and with a learning rate of $\alpha=10^{-3}$ and the exponential decay rates set to $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The quality of the obtained models are assessed in terms of the normalized root mean square (NRMS) and RMS errors:
@@
Fig. 4 shows the simulation performance of the identified models on noiseless test data when trained on noisy data of particular SNRs. Models with various choices of $K_{\theta}$ are also compared: no $K_{\theta}$, linear $K_{\theta}$, and general $K_{\theta}(\hat{z}_{k},u_{k},\hat{e}_{k})$. It can be seen that the linear innovation noise structure is able to reduce the approximation error with a factor of up to two and slight improvements are obtained with the more complex noise model. Next, we analyse the one-step-ahead prediction performance. Similar to the simulation test case, Fig. 5 compares the NRMS errors of these model structures, in a filtering scenario, using noisy test data. As in the simulation scenario, the models are trained on noisy data with the respective SNRs. To provide context for these results, two additional dashed reference lines are included in the figure which represent the one-step-ahead prediction error of the true model, showing the proven convergence and consistency properties as the noise diminishes. It is also clear that incorporating a linear $K_{\theta}$ significantly reduces the NRMS error compared to the baseline model (which is correctly identified) without an innovation noise structure. In overall, both the simulation and prediction results highlight the importance of the innovation noise structure in increasing accuracy of the identified Koopman models.
@@ 5.2 Bouc-Wen benchmark
The Bouc-Wen oscillator benchmark [BW_dataset, NL_benchmark, Schoukens_3_bench] is extensively used for testing capabilities of nonlinear system identification approaches as it describes a system with hysteresis which is a challenging behavior to capture from data accurately. The Bouc-Wen oscillator can be modeled as:
@@
with $m_{L}$ the mass, $y$ its displacement, $\dot{y}$ the velocity, and $u$ the external force applied to the system. The restoring force $r(y,\dot{y})$ is linear while $q(y,\dot{y})$ is a dynamic nonlinear function describing the hysteresis curve. Both these effects are extensively described in [Schoukens_3_bench] where the parameters are chosen for the Bouc-Wen benchmark such that the hysteretic behavior is significantly present.
To estimate a Koopman model of the Bouc-Wen benchmark system, the training dataset contains input-output data generated by using two sinusoidal signals with frequencies of 1 and 4 Hz, as well as four random phase multisine signals with frequency bands ranging from 1 to 150 Hz. For validation, a dataset is generated with an input containing a multisine signal also with an excited frequency range between 1 and 150 Hz followed by a sinusoidal signal at 2 Hz. For test data, the simulated response is generated by multisine and sinesweeep signals as given in [Schoukens_3_bench], as well as a 3 Hz sinusoidal signal to showcase the hysteretic behavior. Note that, as detailed in [Schoukens_3_bench], the data is sampled at a frequency of $f_{\mathrm{s}}=750$ Hz. These used detasets are shown in Fig. 6. As the data is noiseless, the estimation of a noise model is not required which means that $K_{\theta}$ is set to 0.
@@
The best performing model in the multisine case is the model structure with input affine complexity and a lifting dimension of $n_{\mathrm{z}}=100$ (although a close result is also obtained with $n_{\mathrm{z}}=20$). We use this model for comparing the results against other available methods. In Fig. 8, the simulation results using the test data show a low error for the multisine and sinesweep input excitation, and clearly show that the hysteretic behavior is well captured. In Table 1, we can see that the simulation performance of the obtained Koopman model is close to the state of the art (for the interested reader, the other methods are described in more detail in [PLNLSS, Improved_NN_Maarten]). The approaches that obtain slightly better results do not impose a particular structure on the learned model. The Koopman model (25) is able to accurately capture the system behavior and offers a good overall trade-off between state dimension and model complexity. Furthermore, even the more complex structures, i.e., input affine or general, can be cast into LPV representations [Aut_Iacob_inputs] for which there exist convex and computationally efficient control methods [Mohammadpour].
Method gr-$S$-NN SS-NN Suykens Impr SS-NN Suykens Poly-NL-$S$
@@ 5.3.1 Simulation study
The considered simulation model of the drone implements the rigid body dynamics as described in [mahony_multirotor] and uses three coordinate frames: north-east-down (NED) oriented inertial frame $F_{\mathrm{i}}$; the vehicle frame $F_{\mathrm{v}}$ (origin at the centre of gravity of the quadrotor) sharing the same orientation as $F_{\mathrm{i}}$ ; the body frame $F_{\mathrm{b}}$ (orientation fixed to the quadrotor) whose origin coincides with $F_{\mathrm{v}}$. The model has 12 motion states composed of the position s=[xyz]s=[\ x\ \ y\ \ z\ $^{\top}$, translational velocity $v=[\ v_{\mathrm{x}}\ \ v_{\mathrm{y}}\ \ v_{\mathrm{z}}\ ]^{\top}$ expressed in $F_{\mathrm{i}}$, $\zeta=[\ \phi\ \ \theta\ \ \psi\ ]^{\top}$ describing the orientation as Z-Y-X Euler angles in $F_{\mathrm{v}}$, and $\omega=[\ p\ \ q\ \ r]^{\top}$, describing the rotational velocity of $F_{\mathrm{b}}$ w.r.t. $F_{\mathrm{v}}$, given in $F_{\mathrm{b}}$. The inputs to the system are the total thrust $T$ and the torque vector $\tau=[\ \tau_{\phi}\ \ \tau_{\theta}\ \ \tau_{\psi}]^{\top}$ both given in $F_{\mathrm{b}}$ and produced by the four rotors. The diagonal values of the inertia matrix of the drone are set to $J_{\mathrm{x}}=J_{\mathrm{y}}=1.4\times 10^{-5}\,\text{kg}\!\cdot\!\text{m}^{2}$, $J_{\mathrm{z}}=2.17\times 10^{-5}\,\text{kg}\!\cdot\!\text{m}^{2}$, and the off-diagonal values are 0, the mass is $m=0.027\,\text{kg}$, and $g=9.8\,\text{m/s}^{2}$, which are experimentally obtained using a real-world Crazyflie 2.1 quadrotor. The simulation is performed with Runge-Kutta 4 integration at a sampling rate of 48 Hz, accurately replicating the expected flight-dynamics.
As the system is unstable, flight-trajectories are generated by using a gain-scheduled linear quadratic regulator (LQR) controller, designed w.r.t. the local linearisations at each time step of the simulation model. The LQR is scheduled based on a desired state trajectory that is calculated for a $x$-$y$-$z$-$\psi$ defined flight-path reference by taking advantage of the differential flatness property of the system, see [mellinger_control]. We presume full state measurements, hence the collected dataset consists of the system inputs and the states of the dynamical system. Since the dynamics governing the evolution of the position states consist solely of integration, the identification process can be simplified by focusing only on $v^{\top}\ \ \zeta^{\top}\ \ \omega^{\top}\ ]^{\top}$. The size of the recorded datasets can be viewed in Table 10. For more details about the simulation and data collection procedure, the reader can refer to [deep_learning_of_vehicle_dyn].
In our study, we have investigated various dependencies of the input matrix $B$ and emulated different levels of sensor noise. Additionally, we examined the effect of including the original motion states among the observables by setting C=[I 0]C=[\ I\ \ 0\ ] where $I$ and 0 denote identity and zero matrices of appropriate dimensions. This modification aids the design of reference tracking controllers, as it allows the reference signal to be defined directly in terms of the original motion states.
@@
Model performance across various $n_{\mathrm{z}}$ values can be seen in Table 2 in terms of the 160-step NRMS errors for the trained models on the test data set. For comparison purposes, results on a full nonlinear model estimated by the SUBNET approach [deep_learning_of_vehicle_dyn] are also included to indicate performance of a nonlinear model estimate without any restrictions on the network structure. In the network structure column, the subscripts $(\cdot)_{\theta}$, $(\cdot)_{\text{lin}}$, and $(\cdot)_{\mathbf{I}}$ denote the implementation of a function as a deep neural network, linear or identity layers. The superscript of $A^{(\cdot)}$ denotes the lifted state-space dimension. The Koopman model achieves good performance in terms of the NRMS error only slightly exceeding that of the nonlinear SUBNET. Similar to the Bouc-Wen example, we found that the dependence of $B$ on $u$ slightly deteriorates the performance, compared to the input affine model structure. In the table, the effects of measurement noise can also be seen. Even at a level of 20 dB, the Koopman model remains capable of capturing the dynamics. Enforcing the original states among the observables only slightly decreases the simulation precision of the model, which we consider as a good tradeoff for the simpler model structure.
@@
Simulation
$A^{20}$, $B_{\theta}(z)$, $C_{\text{lin}}$ None 0.108 0.120 0.130
@@
Real
$A^{40}$, $B_{\theta}(z)$, $C_{\text{lin}}$ Sensor33footnotemark: 3 0.163 0.213 0.228
@@
66footnotetext: No artificial noise was added, the real dataset was only affected by the noise inherent to the inaccuracies of the various sensors equipped on the Crazyflie.